03:05 UT The 27-m is back in service, so I am grabbing a quick 3C84 observation to check, and possibly adjust, the delays. I will plan to do a two-source observation (3C273 and 3C286) overnight, using 3bands.fsq.
03:50 UT I got good delays and set them, so now I will take a few minutes of 3C84 data as a further test with them applied. The HA on the big dish is now 49.5 degrees, and the HA limit is set to 50 degrees so that the cable alarm does not get set again by going too far. So I should see the dish hit the soft limit in a few minutes. Yep, right on schedule.
04:06 UT I have verified the delays are good (except ant 4 is still showing a lot of noise, so no delay solution). I have now set up my two-source observation to start at 07:00 UT.
10:22 UT So far, the observations have gone flawlessly, so this should be good data to get a baseline update for Bx, By and Bz. I will also be able to check pointing and feed unrotating.
13:30 UT Huge news! The 300 MHz correlator is working. The only known glitch is an off-by-one-slot error for the correlated data, which Jack is working on. Meanwhile, I took data all night on 3C273 and 3C286 using the new/complete science-channel definitions. The data recorded okay, and the science channels are good. However, we had no phase coherence. I discovered that the clock rate was set to 0.8 (GHz) in DPPparameters.f90. I have changed to 1.2, and am taking new data on 3C286 (the stronger 3C273 has set...).
15:00 UT I am back to the 200 MHz correlator, and will take some calibration (two-src) observations on 2136+006 and 2253+161, using 3bands.fsq. The scan started at 15:00 UT, but in fact with the new HA limits the source is not reachable yet. Should be reachable by the next file, at 15:10 UT.
15:30 UT Some sort of strange glitch occurred on Ant 14, and it stopped tracking without a visible error indication. I had to reboot it to recover. It is tracking again now.
18:25 UT I got a nice delay solution on source 2136+006, using delay_widget.py, and updated it. I am now taking some additional data on that source to check the delays. I am reasonably confident of the main (X) delays, but the Y-X delays may have the wrong sign.
19:07 UT I verified that delay_widget used the wrong sign of Y-X delays, so I fixed that and did one more test. The results are great! Just because I am paranoid, I am doing one last test after a very minor update in the delays. Then I will move on to test Jack's latest compile of the 300 MHz correlator.
01:32 UT The 300 MHz correlator appears to be fully working! I have been working with dppxmp all afternoon to solve the phase wind issues, and I believe I have finally solved it. With an 800 MHz clock, our 600-1200 MHz IF was in the third nyquist zone, so the order of channels was forward (low subchannels -> low frequency, etc.). Now with our 1200 MHz clock, the IF is in the second nyquist zone, so the order is reversed. I understood that, and adjusted chan_util_bc.py to write the channel assignments in the inverted order. However, that still did not solve the phase wind. Thinking further, I finally realized that the sign of the phase correction should also reverse. I applied that sign reversal, and now some 3C84 data I just took may be right. Oddly, though, it is only good on baseline 8-14, which is a problem I noted last month, which was only corrected by rebooting the ROACHes. I will do that now...
02:20 UT The phases are great now! We definitely have a working correlator. I adjusted the delays using delay_widget.py, but a couple of baselines required huge delays (of order 10 ns), so I am not sure of the sign. Corrections of +10 and -10 were about equally viable. I'll see what the next scan shows...
05:39 UT In looking at the new 3C84 data after application of the delays, I realized that the delays measured with the new correlator need to be applied with the opposite sign also. I updated delay_widget.py accordingly, and updated the delays with the correct sign, but I have not yet verified the new delays. I have set up to observe 3C273 and 3C286 all night, starting at 06:10 UT.
14:18 UT We are continuing with the 300 MHz correlator, but after running for some hours it tends to start dropping packets and requires a reboot. That creates new delay settings that will affect comparison with calibration data taken at other boot cycles. There are also potential problems with the phase correction (fringe stopping) that have not been fully understood, so data with the new correlator may not be scientifically useful. Still, I believe we need to continue with it to learn more about the failure modes and get them solved. I took 1 h of data on 3C84 last night, but had to reboot both just before and just after, so I am not sure how valuable it will be.
01:23 UT Very strange. The correlator ran perfectly all day. At about 01:05 UT I checked, and it was still good, on the last scan of the day (hour long 3C84 run). After those observations ended, I rechecked and found that the correlator is now dropping packets! How does it know anything about ending observations? It is very puzzling what could affect it in any way. I have rebooted it once again, so it will be interesting to see if it runs all night.
Shaheda reported that IDB files stopped at 14:57:32. Checked packets from dpp and the numbers are 154000-155000. Restarted roaches at 16:30. Still no IDB files. dppxmp is running. Captured some packages and the data are all zero.
18:50 UT Analyzed last night's 3C84 observations and updated delays. Restarting solar scan now (18:54 UT) so they take effect.
23:50 UT Ant 7 was finally repaired with a new jack screw, and is now in operation.
19:14 UT Another 3C84 observation last night shows that the above delay update worked well (except for Ant4, which may have had the wrong delay). A small adjustment in delays was done based on these observation, including a quite small delay for Ant7, and the delays were updated before today's observations were started. A check of the first calibrator scan shows that all delays look pretty good, except Ant4 is still a problem.
12:55 UT The 3C84 observation last night completed successfully, and in addition, I ran the two_src.scd file on 3C273 and 3C286, using 3bands.fsq and alternating sources every 10 min. The 27-m dish tracked correctly during the entire night, so these observations should be suitable for baseline corrections to all three components, . Only Ant12 was not tracking, although the Ant4 delays are probably still off, and may compromise results on that antenna.
13:22 UT The baseline corrections based on the Mar 23 observations were added. Also, I determined a good delay for Ant 4 based on CIEL-2 observations, so I have some hope that it's data will be better. The correction was 48 ns! I have set up to do a 3bands.fsq observation of 2136+006 before sunrise, in order to check the delays at least. The baseline corrections can only be checked by another long observation, so I will plan to run the two_src.scd observation overnight tonight.
19:16 UT This morning's calibration observations revealed that the corrections helped, but were not sufficient. On investigation, I found that the corrections were in m, while the software expects them in ns. Therefore, the correction magnitudes should have been increased by about a factor of three. I have updated the corrections by dividing by mperns (distance in m that light travels in 1 ns), and we will repeat the observation. Unfortunately, it is too windy today to do more.
11:26 UT The solar (and probably calibrator) data show both amplitude and phase oscillations that are not understandable. I rebooted everything except the ROACH boards yesterday, but with no change. Now I am rebooting the ROACHes, in the hopes that something will change. I wanted to avoid it, due to its effect on delays, but there seems no other option.
16:02 UT Updated delays just now, based on this morning's 2136+006 calibration. The delays were relatively large for Ant 14, averaging almost -15 ns, which more or less agrees with a packet capture on CIEL-2 from this morning. Changes are effective as of 16:02 UT.
10:49 UT Natsuha set up for observing multiple sources using 3bands.fsq last night, but continued high winds have compromised the observations. With 20 minutes more to go (two scans), only a few scans had good tracking on the 27-m antenna. The start times of good scans are: 0715, 0805, 0815, 0825, 0855, 0905, 0915, 0925, 1015, and 1025. So far so good on 1055 scan, too. And 1105 is good. Some scans have long slew times, though (2-3 minutes).